Life, the Universe, and Everything

Introduction

Welcome to “Life, the Universe, and Everything”, a panel question and answer event sponsored by the MIT Atheists, Agnostics and Humanists (MITAAH). In this program, you will find summaries of the various religions represented on the panel.

We are very pleased to have on our panel: Brian Aull, representing Baha’i, Nate Osgood, representing Buddhism, John Wuestneck, Connie Parvey, and a third person, not determined at press time representing Christianity, Tom Ferrick, representing Humanism, Imam Taalib, representing Islam, Ilana Goldhaber-Gordon, representing Judaism, and Linda Julien representing Paganism.

Format/Schedule

7:00 - 7:10  Introduction
7:10 - 8:00 Brief descriptions by the panelists
8:00 - 9:00  Moderated Q&A

We will be handing out question forms at the beginning, so that as the panelists are speaking you can write down questions and have them asked.

About MITAAH

MITAAH, the sponsor of this event is a student activity at MIT founded in 1996. More information is available at http://www.mit.edu/˜ mitaah/. Our next meeting will be on Tuesday, April 21st at 7pm in Private Dining Room #3 in the Student Center. If you would like to see future announcments of MITAAH events, feel free to add yourself to the mitaah-announce mailing list on athena by doing:

athena% blanche mitaah-announce -a $USER

Baha’i

Summary by Brian Aull

What is/are the basic tenet(s) of your faith?

Baha’u’llah (1817-1892), the Prophet-Founder of the Baha’i Faith, taught that there is one God Who progressively reveals His will to humanity. Each of the great religions brought by the Messengers of God – Moses, Buddha, Krishna, Jesus, Muhammad, Zoroaster – represents a successive stage in the spiritual development of civilization. Baha’u’llah, the most recent Messenger in this line, has brought teachings that address the moral and spiritual challenges of the modern world. At the heart of Baha’i belief is the conviction that humanity is a single people with a common destiny. In the words of Baha’u’llah, ”The earth is but one country, and mankind its citizens.” Other core teachings include the independent investigation of truth, the elimination of prejudice, the equality of the sexes, the harmony of science and religion, the elimination of extremes of poverty and wealth, and a world commonwealth of nations. The Baha’i writings uphold a very high moral standard, stressing the importance of honesty, trustworthiness, chastity, generosity, and service to others.

Briefly summarize the history of your faith.

The Baha’i Faith arose in the mid-19th century in Iran. A young merchant in Shiraz, titled the Bab (the ”gate”) claimed to be the return of the 12th Imam prophesied in Shi’ah Muslim traditions. The Bab’s mission was twofold: He first announced to the people of His native land His own station as an independent Messenger and He also prepared the way for the coming of another Manifestation (messenger) of God, a Prophet who would announce His station soon after The Bab.

The next six years marked a dramatic increase in both the number of persons who became followers of the Bab and in the energy spent by the Shi’ah clergy of Iran to stamp out this new religion. Eventually 20,000 Babis would be put to death for their beliefs. The Bab Himself was imprisoned and was executed in July of 1850. Many Babis were also imprisoned. Among them was Mirza Husayn-‘Ali, entitled Baha’u’llah (The Glory of God) by The Bab. Imprisoned for several months in 1853 in Tehran and then exiled to Iraq, in the city of Baghdad in 1863 Baha’u’llah announced to the world His station as the One for Whom the Bab had prepared the way. The majority of the Babis accepted Baha’u’llah’s claim and became known as Baha’is (the followers of Baha).

The Baha’i Faith has spread all over the world and has significant communities in more countries than any other major religion except Christianity.

What is the purpose of life?

For Baha’is, the purpose of life is to know and love God, and to progress spiritually. As in most other religions, prayer and meditation are primary tools for spiritual development. However, acts of service to humanity are also essential to spiritual growth, and another purpose of life is to carry forward an ever-advancing civilization. The Baha’i teachings elevate work performed in a spirit of service to the rank of worship. The complementary nature of prayer and action are reflected in the very structure of the Baha’i Houses of Worship, consisting of centers for prayer and mediation surrounded by institutions that carry out charitable and humanitarian functions.

What happens to us after death?

As in the world’s other religions, the Baha’i concept of life after death is deeply integrated into teachings about the nature of the soul and the purpose of this earthly life.

Baha’u’llah confirmed the existence of a separate, rational soul for every human. In this life, He said, the soul is related to the physical body. It provides the underlying animation for the body, and is our real self.

The soul does not die; it endures everlastingly. When the human body dies, the soul is freed from ties with the physical body and the surrounding physical world and begins its progress through the spiritual world. Baha’is understand the spiritual world to be a timeless and placeless extension of our own universe– and not some physically remote or removed place.

Entry into the next life has the potential to bring great joy. Baha’u’llah likened death to the process of birth. He explains: ”The world beyond is as different from this world as this world is different from that of the child while still in the womb of its mother.”

What is the origin of man? Of the universe?

Baha’u’llah teaches that the physical creation is without beginning in time. It is a perpetual emanation from the Creator. Worlds and systems undergo transformation, composition, and decomposition, but their component elements remain. The ”creation” of our universe was not ”making something out of nothing”, but rather was a bringing together of elements that had previously been scattered. Baha’u’llah affirms that the particular universe that gave rise to our planet goes back billions of years. He even described, in terms that seem to hint at a quantum theory of the Big Bang, how this universe came about through the ”heat generated from the interaction between the active force and that which is its recipient.”

Baha’is also see no contradiction between modification of species by natural processes and the notion of creation having a Creator and a purpose. The human species is millions of years old, having evolved from less complex biological forms. However, the Baha’i Faith teaches that man was a distinct species from the beginning, and even when our remote ancestors were simple organisms physically, they possessed a distinct spiritual potentiality.

Where does morality come from? How do we know what is right?

Baha’u’llah teaches that God conferred upon human beings ”the unique distinction and capacity to know Him and to love Him.” So people have the ability to appreciate the beauty of God, as revealed in a pervasive and general way in creation, and in a specific way through His Messengers.

These Manifestations of God – the spiritual luminaries that have founded the great religions of humanity – are the primary moral and spiritual educators of the human race. Through their lives and example, we get our clearest glimpse of the attributes of the Creator. Through their influence, the human soul is awakened. Through the Word of God, revealed by each one of these Beings, ethical standards are established and renewed.

References and more information Books:

The Baha’i Faith: The Emerging Global Religion by William Hatcher and Douglas Martin, Harper & Row, San Francisco, 1984. A text suitable for rigorous study at an introductory level.

The Baha’is: A Profile of the Baha’i Faith and Its Worldwide Community. An in-depth magazine-size publication by the Office of Public Information at the Baha’i International Community. Also available on the oneworld Web site listed below.

Web sites:

http://www.bahai.org/ http://oneworld.wa.com/bahai/magazine/homepage.html http://web.mit.edu/mitba/www/

For a profile of some of the Baha’i community’s race relations efforts, see

”The Power of Race Unity”, which will be aired on Channel 68 WABU on Sundays April 26, May 10, and May 24 at 11 am, and on Fridays May 1, 8, 15, 22 at 8:30 am.

Buddhism

Summary by Nate Osgood.

What is/are the basic tenet(s) of your faith?

Buddhism is best described as an approach to life rather than as a faith (the teachings strongly discourage absolute belief without understanding born of one’s own direct experience ). The basic nature of this viewpoint and approach is as follows: Our experience in this world is imperfect: There are good things we can’t get or that eventually disappear, bad things that we want to avoid but must deal with. We suffer a lot because of this. Some of this suffering is inevitable (e.g. physical pain and bad health), but the large majority of it is quite unnecessary, and arise instead from poor mental health. Most of this unnecessary suffering arises from deeply ingrained mental habits that color how we see the world, how we view ourselves, and how we relate to others - and, ultimately, how we act. Common manifestations of these habits are the types of anxiety and neuroses so common in modern life: Cravings and aversions for material things, hatred or disdain for people different than us, compulsive worrying about our ranking on some judgement scale, anger or sadness about something that happened some time ago, agitated fear of a challenge we will face, etc. Fortunately, there is a way to live less habitually and more consciously. An important step in moving beyond such habits it is to recognize them as products of delusions about the way the world really is: At base, much of our mental suffering rest on refusals to acknowledge that the world is constantly changing, or (even more deeply) on the superstition of an immutable ”self”. But an intellectual understanding of such delusions is not enough: Before we can lessen the grip of these mental habits, we must integrate this realization into our daily life and emotional structure. Meditation serves both to allow us to uncover our delusions and to replace them with a deep understanding that ramifies in both our intellect and emotions. Cultivating such mindfulness awareness serves automatically to lessen our defensiveness and fosters both greater compassion for other conscious creatures and appreciation for the pleasant aspects of life. Good books on this include ”What the Buddha Taught” (Rahula) and ”Seeking the Heart of Wisdom” (Goldstein and Kornfield. Additional Q&A can be found at the BAMIT homepage, at http://www.mit.edu/activities/metta/home.html)

Briefly summarize the history of your faith.

detail) Although it has many roots that extend back far longer, Buddhism first emerged around 600BC in Northern India with the life of Siddartha Guatama. Siddartha abandoned his sheltered life as a prince and took up life as a wandering ascetic, but ultimately proposed a ”middle path” between epicurean indulgence, craving, and self-aggrandizement on the one hand, and ascetism, aversion to life, and desire for self-annihilation on the other. The Buddha’s teaching was iconoclastic, universal and experiential, stressing meditation and mindfulness for all people, regardless of class, sex, or background. Always intent on making a practical difference in people’s day-to-day lives, the Buddha urged his followers to avoid accepting his teaching on faith and to investigate it on their own, and sometimes recommended potential recruits to remain with their existing teachers. Following the Buddha’s death, the story of Buddhism is one of tolerance, local adaptation, and consequent diversity. Buddhism spread first to SE Asia, and later to Central Asia and East Asia as well. The extent of this spread is remarkable in that it was accomplished without coercion or bloodshed. In each of these countries, the tolerance implicit in the approach led to an absorption of many pre-Buddhist practice and beliefs into local Buddhist variants. This led to the eventual delineation of the three major existing approaches (with various schools of closely-related thought in each): Theravada (Sri Lanka and SE Asia), Mahayana (East Asia), Vajrayana (Tibet, Central Asia).

What is the purpose of life?

Buddhism does not view life or human beings as having been created by a God. There is no inherent, God-given ”purpose” to life in the normal religious sense. But the sentient creatures that exist in this world naturally seek happiness and fulfillment, and we can speak loosely of this as a natural ”purpose” (much as we speak of biological organs having ”functions” despite the fact that they were not consciously designed by a purposeful entity). Viewed in this way, our purpose is to help both ourselves and other conscious creatures to live a happy and fulfilled life. (Buddhists view the achievement of genuine happiness for any particular person as requiring a conscious and mindful approach to life – an approach which cannot help but decrease our tendencies towards self-aggrandizement and foster compassion for other beings.)

What happens to us after death?

The Buddhist approach does not see ”death” as a singular event – it is just the final change in a life that is always in flux, always changing. It is merely an illusion to think that there is an unique, well-defined ”I” who has lived a life and now dies. In a very real sense, we are always dying, and always being reborn within each moment of this life. Many Buddhists traditionally believe that this process of birth-death-rebirth continues after the end of life as well. In this view, the ever-changing ”life process” (more akin to an ever-changing flame than an eternal soul) transfers at death to one or more other destinations (new lives). The ”personality” does not live on, but the life process does. In modern times, many other Buddhists (including prominent teachers and monks) argue instead that the processes of ”birth and rebirth” with which we are familiar in each moment are limited in operation to this life. From this perspective, the life process ceases at death. True to its history of as an ”a la carte” religion, Buddhism’s ”big tent” has no difficulty accommodating both viewpoints.

What is the origin of man? Of the universe?

Buddhism takes no particular position taken on the origin of man or of the universe: It is seen as a rather academic question, not really relevant to the much more pressing issue of lessening suffering and finding a way to genuine, compassionate happiness within this life.

Where does morality come from? How do we know what is right?

Buddhism has an ethical code very different from that seen in most other religions. Most notably, Buddhist ethics is teleological rather than deontological. Rather than dichotomizing acts according to some cosmic scale of ”good” and ”evil” or ”right” and ”wrong” a particular act is instead appraised more impassively according to its many consequences - typically some desirable, and some undesirable, some psychological, some physical, some short-term and some longer-term. Though an unskillful act may have myriad undesirable effects (i.e. causing great immediate suffering and perhaps strengthening unskillful mental habits in the mind of the perpetrator or others), it is not viewed as a ”sin”

- in Buddhism, there is no soul to be blemished, and no God to be offended. Buddhist ethics is also distinctive in that it tends to focus not on rule-based reasoning for particular moral situations (something which is seen as overly simplistic and leading to weak resolution), but on self-cultivation of practitioners to the point where they see the effects of situations clearly enough that they are not tempted by unskillful paths, and cannot help but act in moral ways. Morality is thus seen as the most important outward manifestation of general mindfulness; a person in good mental health will act wisely and will not want to be malicious. Finally, the sphere of Buddhist morality embrace not only other human beings, but also all conscious creatures; from a moral standpoint, humans and different species of animals are viewed as differing only in degrees. Some people used to strict, rule-based moral codes may find this contextual approach bewildering and distasteful; to a practicing Buddhist, it simply mirrors the genuine moral complexity of real-world situations.

Good references in this area would be Ken Jones’ ”The Social Face of Buddhism”, and http://jbe.la.psu.edu/1/whitabs1.html

Christianity

Summary compiled by Matthew Gray (due to time commitments of Holy Week, the chaplains were unable to prepare this summary).

General Summary

“For God so loved the world that he gave his only begotten Son, that whoever believes in him shall not perish but have eternal life” (John 3:16)

“There is no man who does not sin” (1 Kings 8:46) “Christ died for sins once for all, the just for the unjust, in order that He might bring us to God” (1 Peter 3:18).

“ ‘I am the way, and the truth, and the life; no one comes to the Father but through Me’ ” (John 14:6)

References and more information Web sites:

http://www.gospelcom.net/ http://www.mit.edu/activities/christro/grouplist.html

Humanism

Summary by Tom Ferrick.

General Summary

Some of the basic tenets of Humanism are that human beings are a natural part of the evolving life in the universe, that we each have a brain (therefore mind) that is commensurate with the reality in which it was formed; knowledge is therefore a survival mechanism. Humanists are monists, disbelieving in two separate worlds of spirit and matter. Mystery abounds but miracles are limited to the imaginative invention of humans. This would include all claims of supernatural revelation. The meaning, purpose, and morality of our lives are what we choose to say they are, philosophical foundations that provide us the surest form of self-identity. Concepts such as personal freedom, democratic justice, the envoronment, and (for art and humor) a sense of proportion are cornerstones of our world-view; there are more, of course. And humanism, being at home with science and history, is organic, every growing and shifting, and never precisely what one person today says it is.

References and more information Web Sites:

http://www.infidels.org is a comprehensive “Yahoo” like listing of secular resources. http://www.mit.edu/˜ mitaah/ is the MITAAH home page and contains our library which includes many good books on Atheism, Agnosticism and Humanism.

Islam

Summary by Seif-Eddeen Fateen.

What are the basic tenets of the Islamic faith?

The first tenet of Islam is the belief in the existance of one and only one God (Allah) who has no partner, parents or children. He is the creator of everything and He was not created. He is all powerful, and He is absolutely just. There is no entity in the entire universe worthy of worship besides Him. He is Eternal, Living, All-Powerful, All-Knowing, All-Seeing, All-Hearing Creator and Sustainer of the Universe.

The second tenet is the belief in Allah’s messengers. Allah has sent messengers to guide all of mankind. These messengers include Moses, Jesus and Muhammad (peace be upon them). All messengers were sent to confirm the belief in one God and Muhammad is the last messanger for all mankind.

The third tenet is the belief in Allah’s books which were revealed to the messangers. Allah’s last book is the Qur’an, which is His own words and is the only authentic revealed book in the world that has been kept without change.

The fourth tenet is the belief in the day of judgement in which every single soul will be held accoutable for his or her beliefs and actions. Righteous people will enter paradise where they will dwell forever whereas those who went astray will enter hell where they will have eternal punishment.

For more information, visit: http://www.introductiontoislam.com/

History of Islam

Because ”Religion with Allah is Islam” (Quran 3.19), Islam started with the first human being, Adam, who taught his offspring how to worship Allah. With time, generations started to deviate from the straight path. Hence, Allah sent many messengers to the people to guide them to Islam. Here are some of the quranic verses that talks about different messengers:

Abraham and his son asked Allah: ”Our Lord, make of us MUSLIMS, bowing to Thy (Will), and of our offspring a nation MUSLIM, bowing to Thy (will)” (Quran 2.128).

”Moses said: ”O my people! If ye do (really) believe in Allah, then in Him put your trust if ye are MUSLIMS” (Quran 10.84).

”When Jesus found Unbelief on their part He said: ”Who will be My helpers to (the work of) Allah?” Said the disciples: ”We are Allah’s helpers: We believe in Allah, and do thou bear witness that we are MUSLIMS” (Quran 3.52)

At the time of Muhammad (pbuh), the earth was full of corruption and idols worshipping. Very few people were following the true message of Islam conveyed by the previous messenger Jesus (pbuh). After the revelation came to Muhammad (pbuh) at the age of 40, he spent 13 years in Mecca trying to invite his people to Islam but very few accepted it. He left to Madina to escape prosecution and to start a new islamic community. In about 80 years, Islam propagated from Madina to cover the whole civilized world at that time, from China in the East to the Atlantic Ocean in the West.

Now Islam is the fastest growing religion in the US and the world with 1.2 billion followers around the globe.

For more information about the life of the prophet Muhammad (pbuh), check out: http://www.erols.com/ibrahimshafi/

What is the purpose of life?

”Blessed is He in Whose hand is the kingdom, and He has power over all things, Who created death and life that He may try you– which of you is best in deeds; and He is the Mighty, the Forgiving.” (Quran 67:1-2). This life is a place of test and the hereafter is a place of reward. ”Surely We have made whatever is on the earth an embellishment for it, so that We may try them (as to) which of them is best in works.” (Quran 18:7)

What happens to us after death?

Death is the transfer from the phase of test to the phase of judgment. ”Every soul shall taste of death, and you shall only be paid fully your reward on the resurrection day; then whoever is removed far away from the fire and is made to enter the garden he indeed has attained the object; and the life of this world is nothing but a provision of vanities.” (Quran 3.185)

What is the origin of man?

God created man from clay and water. Adam was the first man and he descended from Heaven to the earth as a consequence of his disobedience.

The creation of each individual in the womb of his or her mother is mentioned in many verses of the Quran, e.g. ” Then We made the seed a clot, then We made the clot a lump of flesh, then We made (in) the lump of flesh bones, then We clothed the bones with flesh, then We caused it to grow into another creation, so blessed be Allah, the best of the creators.” (Quran 23:14)

What is the origin of the universe?

God created the universe. ”Do not the Unbelievers see that the heavens and the earth were joined together (as one unit of creation), before we clove them asunder? We made from water every living thing. Will they not then believe?” (Quran 21:30)

The origin of the universe was smoke. ”Then turned He to the heaven when it was smoke, and said unto it and unto the earth: Come both of you, willingly or loth. They said: We come, obedient.” (Quran 41:11)

The universe is expanding. ”And the heaven, We built it with might, and most surely We are making expansions” (Quran 51:47)

For more information about science in the quran, check out: http://www.itis-truth.org/

Where does morality come from? How do we know what is right?

Allah is the only source of information regarding what is right and what is wrong. Allah created us from nothing; His vast knowledge, wisdom and power can not be matched by those of humans, and He is the best to know what is good and what is bad for His creation. Thus, in defining what is right and what is wrong, Muslims follow the general guidelines they learn from the Quran and the words and actions of the prophet Muhammad (pbuh) which are called Sunnah. Quran and Sunnah provide a significant level of details as to how Muslims conduct their lives and perceive the world around them. Muslims also have the responsibility of working their minds in understanding the Quran and Sunnah and extending that understanding to face new challenges. Nevertheless, Muslims realize that the basic guidelines of morality, as defined by Allah, are absolute regardless of time and place. Every action is classified as right or wrong within the framework of Quran and Sunnah.

References and more information

For general information about Islam, check out the following sites: http://www.islamic.org.uk/ http://www.usc.edu/dept/MSA/ http://www.islamworld.net/

Questions? Don’t hesitate to contact: Abdullah Taalib Faruuq Imam of Masjid Lilhamdulellah 442-2805

Suheil Laher MIT Muslim Chaplain slaher@ptechinc.com

Seif-Eddeen Fateen President, MIT Muslim Students Association sfateen@mit.edu, x2-1502

Judaism

General Summary

Summary by Ilana Goldhaber-Gordon.

Judaism is a religion of actions. We believe that God judges us on what we do, not on what we think or believe. We also believe that by repeating actions that demonstrate a commitment to religion, God and community, a person will come to actually feel that commitment. In other words, actions are important because they define who we are, both in our own minds, in other people’s eyes, and in God’s eyes. Worthy actions fall into two general categories: good deeds that help other people, and ritual actions that may seem arbitrary and are intended to demonstrate obedience to the word of God. Together, these two categories touch every human action: eating, praying, marrying, dying, using the bathroom, making love, dressing, undressing, going to sleep. We try to infuse every aspect of our lives with a touch of divinity, and so we combine rituals and prayers with almost everything we do. In addition, our actions connect us to a chain of generations that, we believe, have been performing the same rituals for four thousands of years.

Second only to action is education; according to some rabbinic texts, education even supersedes action. Judaism considers the study of “Torah”, Jewish law and philosophy, to be the most valuable activity. Education must include questioning, and it is best for two students sit together over text, working out it’s meeting and debating it’s ideas between themselves. Teachers should offer occasional guidance, but the real learning is up to the students.

Today, most Jews identify with one of three branches of Judaism: Orthodox, Conservative, and Reform. We Orthodox Jews think of ourselves as the keepers of tradition. We aspire to live according to the principles and traditions of our parents and their parents. A difficulty with Orthodox Judaism is that the tradition provides no means of change. The Rabbis do not have the authority to change the law (Rabbis are simply learned individuals who can provide guidance and teaching), and so changes must occur by a very gradual and subtle evolution. Conservative Judaism also aspires to maintain the traditional lifestyle, with one exception.

Conservative Judaism sanctions it’s Rabbis to make, when necessary, particular changes in Jewish law or custom. Reform Judaism represents a far more major change in philosophy, considering religious law to be a matter of personal preference. According to Reform philosophy (as best as I can represent it) a Jew should focus first on being a good person, and should use the traditions and rituals as he or she sees fit.